Given f (x2) =x2sinx², compute the derivative of f with respect to x.


1
Expert's answer
2020-04-30T18:58:53-0400

As "f(x^2)=x^2sinx^2"

"\frac{d(f(x^2))}{dx}=\frac{d(x^2sinx^2)}{dx}"

"2xf'(x^2)=2xsinx^2+2x^3cosx^2"

Hence, "f'(x^2)=sinx^2+x^2cosx^2"